OpenStudy (anonymous):

An Olympic runner can run 5000 meters in 12 minutes. How long would it take the runner to finish a 42,000 meter run?

OpenStudy (valpey):

Assuming the runner doesn't tire, are you kidding me?

OpenStudy (anonymous):

since v=s/t therefore : v=5000/(12X60) .... so ..... v=7m/s therefore the required time (t) = s/v = 42000/7=6000 sec. Note: this answer is approximated

OpenStudy (valpey):

42,000 m is 42/5 as far as the 5000 m run so it should take 42/5 times as long (assuming the runner is a machine).

OpenStudy (anonymous):

its minutes. would it be 100.8? 42000/5000=8.4*12=100.8

OpenStudy (lgbasallote):

that olympic runner is that runner in temple run

OpenStudy (lgbasallote):

though i dont know if anyone has achieved 42,000 meters in temple run

OpenStudy (lgbasallote):

@Mathrocks4me 100.8 seems right to me

OpenStudy (valpey):

I get 8.4*12 = 100.8 min or 1 hr 40 min 48 sec.

OpenStudy (anonymous):

yeah ur Answer is right @Mathrocks4me i produced my answer in seconds

OpenStudy (lgbasallote):

you can just do \[\frac{42, 000}{5, 000} \times 12\] though...same result